OpenStudy (anonymous):

Helpppppp !!! Prove that 1-cos² A / 1+sin A = sin A

OpenStudy (raden):

hmmm... that's looks not right. if we use A = 90 degrees, get 1-cos² A / 1+sin A = sin A 1-cos² 90 / 1+sin 90 = sin 90 1- 0 / 1+1 = 1 1/2 = 1 LHS not equal RHS

OpenStudy (anonymous):

thankyou :)

OpenStudy (triciaal):

its right